OpenStudy (inexilewetrust):

Check answer real quick? What is an object's mass if its speed is .8 m/s and its momentum is 156 kg* m/s ? Momentum is the product of Mass and Velocity. 156 kg*m/s (Momentum) / 0.8 m/s (Velocity) My answer is 195 kg.

OpenStudy (matt101):

Right!

OpenStudy (abhisar):

As it says, \(\tt Momentum (P) = Mass(m) \times Velocity(v)\) Momentum is given 156 and velocity is given 0.8m/s. Plugging the values into the equation we get, \(\tt 156=m\times 0.8 \\ \Rightarrow m=\frac{156}{0.8}=?\)

OpenStudy (inexilewetrust):

@abshisar How did you get from 0.8 to 8?

OpenStudy (abhisar):

I thought the speed is given 8 but it's actually 0.8. Check again I have corrected it.

OpenStudy (inexilewetrust):

Okay so it is 195 kg?

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (inexilewetrust):

Awesome, thanks! @Abshihar and @matt101

OpenStudy (inexilewetrust):

*Abhisar Sorry, I am not wearing my glasses!

OpenStudy (abhisar):

Welcome (-: